Grade 12Differential Calculus

differentiate the following function with respect to x
x^(tan^-1(x))

taniska
12 Years agoGrade 12

1 Answer

Rinkoo Gupta
12 Years ago

let y=x^tan-1x

taking log on both sides, we get

logy=tan-1xlogx

Now diif. w.r.to x by using product rule,we get

(1/y). dy/dx=tan-1x.(1/x)+logx.[1/(1+x2)]

dy/dx= x^tan-1x[tan-1x.(1/x)+logx.{1/(1+x2)}]

Thanks & Regards

Rinkoo Gupta

AskIITians Faculty